Project

General

Profile

NDOS Commissioning Meeting 2011/08/01

Attendance: Jon, Athans, Mark Mathis, Martin, Tomas, Kurt, Andrew, Carl B., Leon, Peter

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • NuMI beam returned Friday late afternoon. As of this morning, beam was running smoothly at ~35e12 POT/spill.
  • APD reinstallation plan: at the very earliest we will reinstall APDs when Mat gets back in two weeks. However, tests at CalTech may not be complete by then.

DAQ Status & Plans

  • Kurt: would like to do some tests of CfgMgr; if changes are successful, we will leave the changes in. The changes are the last three connection configuration types to go into the dB. Historical configs will also be uploaded into the dB.
  • Jon: TDU wig-out study? Andrew: in progress, we're monitoring by beating two TDUs against each other. Second TDU was rebooted over the weekend so we lost the comparison buffer. Monitoring code is in progress. Engineers are also looking at the TDUs.
  • Leon:
    • should have ability long-term Bill Gilbert was hopefu
    • tested dummy coated APDs; at least on appears to be ~fully functional. That one may go on the detector for further studies.
    • remnant TECC noise studies:
      • cap-boards do a lot to kill noise, but not complete
      • barite beads on TECC lead seem to kill the remnant noise on 4 FEBs. Ordering parts to see results on another 50 FEBs. Parts should be available soon (days) for installation. Leon will be at FNAL after Lehman review next week and will work with Denis on installation. Should take less than a couple of days.
    • AC on env. chamber doesn't work, so this needs to be fixed before o-ring pressure tests at CalTech can begin.

DCS Status & Plans

  • Athans: progress on IOC is stalled while firmware work is being done. Developments in GUI are in progress. Alarming discussion to continue.

AOB

  • None.